# Twists of Superconformal Algebras

### v2

@@ Abstract.
We develop the theory of “conformal twists” of superconformal field theories in dimension 3 to 6, extending the well-known analysis of twists for supersymmetric theories. The conformal twists describe all possible inequivalent choices of a nilpotent element in the superconformal algebra. Such twists can give rise to interesting subalgebras and protected sectors of operators, with the Donaldson–Witten topological field theory and the vertex operator algebras of 4-dimensional $\mathcal{N}{=}\hskip 1.00006pt2$ SCFTs being prominent examples. We work mostly with the complexified superconformal algebras, unless explicitly stated otherwise; real forms of the superconformal algebra may have important physical implications, but we only discuss these subtleties in a few special cases. To obtain mathematical precision, we explain how to extract vertex algebras and $\mathbb{E}_{n}$ algebras from a twisted superconformal field theory using factorization algebras.
@@ 1. Introduction
The idea of twisting a supersymmetric theory goes back to Witten [Witten_1988], who introduced it as a recipe for constructing topological quantum field theories, and it has provided a number of examples with rich consequences in physics and mathematics. We wish to emphasize a more abstract formulation of the twisting procedure: given a field theory with an action of a super Lie algebra $\mathfrak{g}$ and an odd element $\mathcal{Q}\in\mathfrak{g}$ satisfying $[\mathcal{Q},\mathcal{Q}]=0$, there is a deformation of the theory by the action of $\mathcal{Q}$. For example, if $H_{\mathcal{Q}}$ is the Hamiltonian associated to $\mathcal{Q}$, we may replace the algebra of local operators $\mathrm{Op}_{\mathrm{loc}}$ by its $\mathcal{Q}$-cohomology
@@
where the condition $[\mathcal{Q},\mathcal{Q}]=0$ ensures that the differential does indeed square to zero. For us the pertinent super Lie algebra $\mathfrak{g}$ will be a superconformal algebra (i.e., the relevant spacetime symmetries for a superconformal theory).
In the last two decades, representation theorists have analyzed the space of square-zero odd elements of super Lie algebras (see especially the work of Jenkins and Nakano [JenkinsNakano]). Of particular relevance to us is the work of Duflo and Serganova [DufloSerganova] (see also [GHSS]), who studied simple super Lie algebras over the complex numbers and described the adjoint orbits of such square-zero odd elements of such algebras. Many superconformal algebras, after complexification, are simple, as we discuss in Section 2. For example, for spacetime dimensions three and six, the superconformal algebras correspond to the super Lie algebras $\mathrm{osp}(n|m)$, which were studied in detail by Gruson [Gruson] prior to Duflo and Serganova. In this paper we explain how these structural mathematical results interface with the role of these super Lie algebras as symmetries of superconformal field theories.
The more familiar case of supersymmetry (i.e. of a supersymmetric extension of the group of isometries rather than of conformal transformations) motivates our approach. As mentioned, Witten introduced twisting to obtain a topological quantum field theory out of a supersymmetric theory: if the stress-energy tensor $T_{\mu\nu}$ of a supersymmetric theories is exact for the action of $\mathcal{Q}$, then this stress-energy tensor vanishes in the $\mathcal{Q}$-cohomology of observables  [^1] Note that such theories are sometimes called “Witten-type” topological theories, as opposed to “Schwarz-type” topological theories where $T_{\mu\nu}$ is exactly zero, as for instance in [BBRT]., and the associated twisted theory is topological. Thus, it is natural to study topological twisting in two stages:
@@
Recently, in [ElliottSafronov, ElliottSafronovWilliams] the first author, together with Safronov and Williams, systematically described the topological twists of super Yang–Mills theories (formulated in the Batalin–Vilkovisky formalism) and identified the twists in super Poincaré algebras that would produce topological theories. Note that there are twists that lead to non-topological theories, where subtle structures emerge, such as theories that look holomorphic [Johansen, Baulieu] or look topological in some directions and holomorphic in others [KapustinHolo, KallenZabzine, ACMV]. In the present work our goal is to pursue analogues for superconformal field theories by working with the superconformal algebras that extend super Poincaré algebras.
Working in the superconformal case is not simply a matter of reproducing the calculations from earlier work in yet another example. There are substantially different features that allow us to construct interesting and novel algebras of operators upon twisting, and superconformal transformations produce fascinating relationships between different twists. An important motivation for us — demonstrating some novelties not apparent in the supersymmetric case — is the work of Beem, Lemos, Liendo, Peelaers, Rastelli and van Rees [Beem_2015], who showed how such superconformal twists lead to novel constructions of vertex algebras. Combinations of these authors wrote a series of subsequent papers that address more examples and explore their consequences [Beem_2015b, Beem_2015c, Beem_2017, Beem_2018]. In this paper we do not focus on specific theories, but rather analyze formal aspects of superconformal actions on algebras of observables and the structures that may arise upon twisting. Our paper might be seen as trying to provide some systematic explanation for the kinds of results found by Beem et al. In Section 3 we provide an overview with a physical flavor; in Section 8 we state precise mathematical results using the language of (pre)factorization algebras. (We discuss again the results of Beem et al in Section 5.3.)
@@ 1.1. What Twisting Does to Superconformal Theories
As a superconformal theory, the superconformal group acts equivariantly: for every $g\in\mathrm{SConf}(p,q|S)$ and open set $U$, there is an isomorphism $\mathrm{Obs}(U)\to\mathrm{Obs}(gU)$. Similarly, the super Lie algebra $\mathfrak{sconf}(p,q|S)$ acts by derivations on every $\mathrm{Obs}(U)$. As the observables form a complex vector space, this action extends to an action by the complexified super Lie algebra, which we will denote $\mathfrak{sconf}(n|S_{\mathbb{C}},\mathbb{C})$ for now.
If $\mathcal{Q}$ is an odd element of the complexified superconformal Lie algebra such that $[\mathcal{Q},\mathcal{Q}]=0$, we may study the algebra of observables $\mathrm{Obs}^{\mathcal{Q}}$ in the theory twisted by $\mathcal{Q}$. This twisted algebra of observables still lives over $C(\mathbb{R}^{p,q})$, and a goal in this paper is to develop a systematic understanding of the properties of this twisted algebra determined purely by the superconformal algebra. To this end, let us describe a few natural structures determined by $\mathcal{Q}$ and independent of the choice of theory and ruminate on what they mean for the twisted theory.
@@
    $\mathfrak{z}_{\mathcal{Q}}^{\mathbb{C}}=\{A\in\mathfrak{sconf}(n|S_{\mathbb{C}},\mathbb{C})_{0}\colon[A,\mathcal{Q}]=0\},$
    where $\mathfrak{sconf}(n|S_{\mathbb{C}},\mathbb{C})_{0}$ denotes the even part of the superconformal algebra. The Lie subalgebra $\mathfrak{z}_{\mathcal{Q}}^{\mathbb{C}}$ consists of even symmetries that commute with the supercharge $\mathcal{Q}$. These symmetries remain as symmetries of the twisted theory $\mathrm{Obs}^{Q}$. The action of non-closed symmetries is broken when we pass to the twist.
@@
    $Z_{\mathcal{Q}}^{\mathbb{C}}\subseteq\mathrm{SConf}(n|S_{\mathbb{C}},\mathbb{C})_{0}$
    denote the Lie group exponentiating $\mathfrak{z}_{\mathcal{Q}}^{\mathbb{C}}$, and let $Z_{\mathcal{Q}}$ denote its intersection with the real form $\mathrm{SConf}(p,q|S_{\mathbb{C}})_{0}$. This Lie subgroup consists of the even symmetries that continue to act on the twisted theory $\mathrm{Obs}^{\mathcal{Q}}$ at the group level, by genuine conformal transformations.
@@
    $\mathfrak{b}_{\mathcal{Q}}^{\mathbb{C}}=\{A\in\mathfrak{sconf}(n|S_{\mathbb{C}},\mathbb{C})_{0}\colon A=[\mathcal{Q},\mathcal{Q}^{\prime}]\text{ for some }\mathcal{Q}^{\prime}\}.$
@@
    $B_{\mathcal{Q}}^{\mathbb{C}}\subseteq\mathrm{SConf}(n|S_{\mathbb{C}},\mathbb{C})_{0}$
    denote the Lie group exponentiating $\mathfrak{b}_{\mathcal{Q}}^{\mathbb{C}}$, and let $B_{\mathcal{Q}}$ denote its intersection with the real form $\mathrm{SConf}(p,q|S_{\mathbb{C}})_{0}$. This Lie subgroup provides the even symmetries of the twisted theory that act on spacetime by genuine conformal transformations but are trivialized on the twisted theory.
@@
Suppose that there is a subspace $\mathbb{R}^{2}\subseteq C(\mathbb{R}^{p,q})$ so that the full group of isometries of $\mathbb{R}^{2}$ lies in $Z_{\mathcal{Q}}$, but only the complexified translation $\overline{d}_{z}$ lies in $\mathfrak{b}^{\mathbb{C}}_{\mathcal{Q}}$, where $z$ denots a complex coordinate on $\mathbb{R}^{2}\cong\mathbb{C}$. Then if we restrict observables of $\mathrm{Obs}^{Q}$ to those supported along this subspace, we have an isometry action so that the complexified antiholomorphic translation is trivialized up to homotopy. In this case (and under some additional technical assumptions), the restricted observables can be equipped with the structure of a vertex algebra. This construction was originally considered by Beem et al [Beem_2015] in their derivation of vertex algebras from 4d $\mathcal{N}=2$ superconformal field theory.
@@ Acknowledgements
We thank Philip Argyres, Ingmar Saberi, Mitch Weaver, and Brian Williams for helpful discussions on this topic, and particularly Richard Eager, who offered detailed comments on the manuscript, especially in guiding us through the literature. The National Science Foundation supported OG through DMS Grant 2042052. ML is supported in part by the National Research Foundation of Korea (NRF) Grant 2021R1A2C2012350. CE and OG thank their REU students at Amherst College (Osha Jones and Ziji Zhou) and UMass Amherst (Nikos Orginos, Nathaniel Reid, and Sam Silver) for work on two- and three-dimensional versions of these problems, which gave helpful complementary insights. OG completed revisions on this paper at the Max Planck Institute of Mathematics during a lovely and productive visit there, and he acknowledges its support with gratitude.
@@ 2. Superconformal Algebras and Twists
We rapidly review the conformal and superconformal groups and their associated (super) Lie algebras. For the conformal situation, we follow Schottenloher’s conventions (see Chapter 2 of [Schottenloher]).
@@ 2.1. The Conformal Case
$\iota(v_{1},\ldots,v_{p+q})=\left(\frac{1-|v|}{2},v_{1},\ldots,v_{p+q},\frac{1+|v|}{2}\right).$
@@
$\left(\frac{1-|v|}{2}:v_{1}:\ldots:v_{p+q}:\frac{1+|v|}{2}\right)\in\mathbb{RP}^{p+q+1},$
@@
In fact, these subgroups collectively generate the full conformal group; in other words, we can always factor a conformal transformation as some composition of transformations of those four types. We refer to [Schottenloher, §2.2] and [dFMS, §4.1] for a thorough discussion  [^2] Note that the action of the special conformal transformations on a point in $\mathbb{R}^{p,q}$ can be interpreted as an inversion (about the origin), followed by a translation, followed by another inversion.
@@ 2.1.1. A more physical formulation
We use the customary notation from the physics literature where repeated indices are summed over, i.e. $\partial^{\mu}x_{\mu}=\sum_{\mu=1}^{d=p+q}\partial^{\mu}x_{\mu}$ Furthermore, indices are “raised” and “lowered” by contracting with the metric (tensor) $g_{\mu\nu}$ or its inverse, i.e.: $\partial_{\mu}=g_{\mu\nu}\partial^{\nu}=\sum_{\nu=1}^{d}g_{\mu\nu}\partial^{\nu}$. The metric is given by: $g_{\mu\nu}=\text{diag}(\mathbb{1}_{p},-\mathbb{1}_{q})$.
@@ 2.2. The Superconformal Case
We will focus on superconformal Lie algebras, as they play the key role in this paper and because they are simpler to describe in technical terms (In particular, we will not discuss super analogs of the conformal compactification.) Our approach is borrowed from Shnider’s elegant and efficient discussion [Shnider]. We begin by recalling how one formulates the super versions of the spacetimes $\mathbb{R}^{p,q}$ and of their isometries. That is, we review what super-Minkowski space is and what the super Poincaré group is (in arbitrary signature) to elucidate the logic that motivates a formulation of superconformal algebras.
@@
In looking for a super version of the conformal algebra, we want to construct a super Lie algebra $\mathfrak{sconf}(p,q|S)$ that depends on this choice of a Majorana spinor representation $S$ and the bilinear form. It should include the super Poincaré Lie algebra as a subalgebra, just as the Poincaré Lie algebra is a subalgebra of $\mathfrak{conf}(p,q)$. Moreover, the reduced Lie algebra of $\mathfrak{sconf}(p,q|S)$ should be $\mathfrak{conf}(p,q)$. In other words, $\mathfrak{sconf}(p,q|S)$ should be a common generalization of usual supersymmetry algebra and usual conformal algebra.
@@
When studying super Lie algebras $\mathfrak{g}$, we will use the notation [,][-,-][ - , - ] for the full graded skew-symmetric Lie bracket
$[-,-]\colon\wedge^{2}(\mathfrak{g})\to\mathfrak{g}.$
@@
The remarkable theorem is that such a super Lie algebra exists only when $p+q\leq 6$, as discovered by Nahm [Nahm], although we emphasize here the clean mathematical approach of Shnider [Shnider], who articulates these properties as explicit hypotheses. Shnider begins by studying the complexification of the situation above, since any such real super Lie algebras has a complexifiction and there is no dependence on a choice of signature after complexifying. He then shows that if such a complex super Lie algebra $\mathfrak{g}$ exists, there is a simple super Lie algebra, satisfying the same properties, which is a quotient or subalgebra of $\mathfrak{g}$. Moreover, its bosonic part splits as the sum of $\mathfrak{so}(p+q,\mathbb{C})$ and a complementary ideal. He then uses Kac’s classification of simple super Lie algebras to show that such complex super Lie algebras do not exist above dimension 6. Identifying the relevant simple super Lie algebras in dimensions less than or equal to 6 is possible and already accomplished.
@@
The classification of simple super Lie algebras has played an essential role in the physics literature on the structure and properties of superconformal theories [Mack, Minwalla, BHP]. In particular, Dolan and Osborn [DolanOsborn] developed the representation theory of 4d superconformal algebras, and Córdova, Dumitrescu and Intriligator [CDI] provided a thorough analysis of the representation theory of superconformal algebras in dimensions 2 to 6 and its applications to superconformal field theory.

### v3

@@ Abstract.
We take first steps toward a theory of “conformal twists” for superconformal field theories in dimension 3 to 6, extending the well-known analysis of twists for supersymmetric theories. A conformal twist is a square-zero odd element in the superconformal Lie algebra, and we classify all twists and describe their orbits under the adjoint action of the superconformal group. We work mostly with the complexified superconformal algebras, unless explicitly stated otherwise; real forms of the superconformal algebra may have important physical implications, but we only discuss these subtleties in a few special cases. Conformal twists can give rise to interesting subalgebras and protected sectors of operators in a superconformal field theory, with the Donaldson–Witten topological field theory and the vertex operator algebras of 4-dimensional $\mathcal{N}{=}\hskip 1.00006pt2$ SCFTs being prominent examples. To obtain mathematical precision, we explain how to extract vertex algebras and $\mathbb{E}_{n}$ algebras from a twisted superconformal field theory using factorization algebras.
Key words and phrases:
Super Lie algebra, conformal geometry, superconformal field theory, nilpotence variety, associated variety, vertex algebra, factorization algebra
@@ 1. Introduction
The idea of twisting a supersymmetric theory goes back to Witten [45], who introduced it as a recipe for constructing topological quantum field theories, and it has provided a number of examples with rich consequences in physics and mathematics. We wish to emphasize a more abstract formulation of the twisting procedure: given a field theory with an action of a super Lie algebra $\mathfrak{g}$ and an odd element $\mathcal{Q}\in\mathfrak{g}$ satisfying $\mathcal{Q},\mathcal{Q}]=0$, there is a deformation of the theory by the action of $\mathcal{Q}$. For example, if $H_{\mathcal{Q}}$ is the Hamiltonian associated to $\mathcal{Q}$, we may replace the algebra of local operators $\mathrm{Op}_{\mathrm{loc}}$ by its $\mathcal{Q}$-cohomology
@@
where the condition $\mathcal{Q},\mathcal{Q}]=0$ ensures that the differential does indeed square to zero. For us the pertinent super Lie algebra $\mathfrak{g}$ will be a superconformal algebra (i.e., the relevant spacetime symmetries for a superconformal theory).
In the last two decades, representation theorists have analyzed the space of square-zero odd elements of super Lie algebras (see especially the work of Jenkins and Nakano [29]). Of particular relevance to us is the work of Duflo and Serganova [20] (see also [27]), who studied simple super Lie algebras over the complex numbers and described the adjoint orbits of such square-zero odd elements of such algebras. Many superconformal algebras, after complexification, are simple, as we discuss in Section 2. For example, for spacetime dimensions three and six, the superconformal algebras correspond to the super Lie algebras $\mathrm{osp}(n|m)$, which were studied in detail by Gruson [28] prior to Duflo and Serganova. In this paper we explain how these structural mathematical results interface with the role of these super Lie algebras as symmetries of superconformal field theories.
The more familiar case of supersymmetry (i.e. of a supersymmetric extension of the group of isometries rather than of conformal transformations) motivates our approach. As mentioned, Witten introduced twisting to obtain a topological quantum field theory out of a supersymmetric theory: if the stress-energy tensor $T_{\mu\nu}$ of a supersymmetric theories is exact for the action of $\mathcal{Q}$, then this stress-energy tensor vanishes in the $\mathcal{Q}$-cohomology of observables  [^1] Note that such theories are sometimes called “Witten-type” topological theories, as opposed to “Schwarz-type” topological theories where $T_{\mu\nu}$ is exactly zero, as for instance in [11]., and the associated twisted theory is topological. Thus, it is natural to study topological twisting in two stages:
@@
Recently, the twists in super Poincaré algebras that would produce topological theories were all identified and the topological twists of super Yang–Mills theories (formulated in the Batalin–Vilkovisky formalism) were described in [23, 22]. Note that there are twists that lead to non-topological theories, where subtle structures emerge, such as theories that look holomorphic [30, 5] or look topological in some directions and holomorphic in others [33, 31, 1]. In the present work our goal is to pursue analogues for superconformal field theories by working with the superconformal algebras that extend super Poincaré algebras.
Working in the superconformal case is not simply a matter of reproducing the calculations from earlier work in yet another example. There are substantially different features that allow us to construct interesting and novel algebras of operators upon twisting, and superconformal transformations produce fascinating relationships between different twists. An important motivation for us — demonstrating some novelties not apparent in the supersymmetric case — is the work of Beem, Lemos, Liendo, Peelaers, Rastelli and van Rees [6], who showed how such superconformal twists lead to novel constructions of vertex algebras. Combinations of these authors wrote a series of subsequent papers that address more examples and explore their consequences [7, 9, 8, 10]. In this paper we do not focus on specific theories, but rather analyze formal aspects of superconformal actions on algebras of observables and the structures that may arise upon twisting. Our paper might be seen as trying to provide some systematic explanation for the kinds of results found by Beem et al. In Section 3 we provide an overview with a physical flavor; in Section 8 we state precise mathematical results using the language of (pre)factorization algebras. (We discuss again the results of Beem et al in Section 5.3.)
@@ 1.1. What Twisting Does to Superconformal Theories
As a superconformal theory, the superconformal group acts equivariantly: for every $g\in\mathrm{SConf}(p,q|S)$ and open set $U$, there is an isomorphism $\mathrm{Obs}(U)\to\mathrm{Obs}(gU)$. Similarly, the super Lie algebra $\mathfrak{sconf}(p,q|S)$ acts by derivations on every $\mathrm{Obs}(U)$. As the observables form a complex vector space, this action extends to an action by the complexified super Lie algebra, which we will denote $\mathfrak{sconf}(n|S,\mathbb{C})$ for now.
If $\mathcal{Q}$ is an odd element of the complexified superconformal Lie algebra such that $\mathcal{Q},\mathcal{Q}]=0$, we may study the algebra of observables $\mathrm{Obs}^{\mathcal{Q}}$ in the theory twisted by $\mathcal{Q}$. This twisted algebra of observables still lives over $C(\mathbb{R}^{p,q})$, and a goal in this paper is to develop a systematic understanding of the properties of this twisted algebra determined purely by the superconformal algebra. To this end, let us describe a few natural structures determined by $\mathcal{Q}$ and independent of the choice of theory and ruminate on what they mean for the twisted theory.
@@
    $\mathfrak{z}_{\mathcal{Q}}^{\mathbb{C}}=\{A\in\mathfrak{sconf}(n|S,\mathbb{C})_{0}\colon[A,\mathcal{Q}]=0\},$
    where $\mathfrak{sconf}(n|S,\mathbb{C})_{0}$ denotes the even part of the superconformal algebra. The Lie subalgebra $\mathfrak{z}_{\mathcal{Q}}^{\mathbb{C}}$ consists of even symmetries that commute with the supercharge $\mathcal{Q}$. These symmetries remain as symmetries of the twisted theory $\mathrm{Obs}^{Q}$. The action of non-closed symmetries is broken when we pass to the twist.
@@
    $Z_{\mathcal{Q}}^{\mathbb{C}}\subseteq\mathrm{SConf}(n|S,\mathbb{C})_{0}$
    denote the Lie group exponentiating $\mathfrak{z}_{\mathcal{Q}}^{\mathbb{C}}$, and let $Z_{\mathcal{Q}}$ denote its intersection with the real form $\mathrm{SConf}(p,q|S)_{0}$. This Lie subgroup consists of the even symmetries that continue to act on the twisted theory $\mathrm{Obs}^{\mathcal{Q}}$ at the group level, by genuine conformal transformations.
@@
    $\mathfrak{b}_{\mathcal{Q}}^{\mathbb{C}}=\{A\in\mathfrak{sconf}(n|S,\mathbb{C})_{0}\colon A=[\mathcal{Q},\mathcal{Q}^{\prime}]\text{ for some }\mathcal{Q}^{\prime}\}.$
@@
    $B_{\mathcal{Q}}^{\mathbb{C}}\subseteq\mathrm{SConf}(n|S,\mathbb{C})_{0}$
    denote the Lie group exponentiating $\mathfrak{b}_{\mathcal{Q}}^{\mathbb{C}}$, and let $B_{\mathcal{Q}}$ denote its intersection with the real form $\mathrm{SConf}(p,q|S)_{0}$. This Lie subgroup provides the even symmetries of the twisted theory that act on spacetime by genuine conformal transformations but are trivialized on the twisted theory.
@@
Suppose that there is a subspace $\mathbb{R}^{2}\subseteq C(\mathbb{R}^{p,q})$ so that the full group of isometries of $\mathbb{R}^{2}$ lies in $Z_{\mathcal{Q}}$, but only the complexified translation $\overline{d}_{z}$ lies in $\mathfrak{b}^{\mathbb{C}}_{\mathcal{Q}}$, where $z$ denots a complex coordinate on $\mathbb{R}^{2}\cong\mathbb{C}$. Then if we restrict observables of $\mathrm{Obs}^{Q}$ to those supported along this subspace, we have an isometry action so that the complexified antiholomorphic translation is trivialized up to homotopy. In this case (and under some additional technical assumptions), the restricted observables can be equipped with the structure of a vertex algebra. This construction was originally considered by Beem et al [6] in their derivation of vertex algebras from 4d $\mathcal{N}=2$ superconformal field theory.
@@ Acknowledgements
We thank Philip Argyres, Mitch Weaver, and Brian Williams for helpful discussions on this topic, and particularly Richard Eager, who offered detailed comments on the manuscript, especially in guiding us through the literature. We also thank an insightful, patient, thorough referee whose feedback substantially improved the paper. The National Science Foundation supported O.G. through DMS Grant 2042052. ML is supported in part by the National Research Foundation of Korea (NRF) Grant 2021R1A2C2012350. C.E. and O.G. thank their REU students at Amherst College (Osha Jones and Ziji Zhou) and UMass Amherst (Nikos Orginos, Nathaniel Reid, and Sam Silver) for work on two- and three-dimensional versions of these problems, which gave helpful complementary insights.
@@ 2. Superconformal Algebras and Twists
We rapidly review the conformal and superconformal groups and their associated (super) Lie algebras. For the conformal situation, we follow Schottenloher’s conventions (see Chapter 2 of [42]).
@@ 2.1. The Conformal Case
$\iota(v_{1},\ldots,v_{p+q})=\left(\frac{1-|v|^{2}}{2},v_{1},\ldots,v_{p+q},\frac{1+|v|^{2}}{2}\right).$
@@
$\left(\frac{1-|v|^{2}}{2}:v_{1}:\ldots:v_{p+q}:\frac{1+|v|^{2}}{2}\right)\in\mathbb{RP}^{p+q+1},$
@@
In fact, these subgroups collectively generate the full conformal group; in other words, we can always factor a conformal transformation as some composition of transformations of those four types. We refer to [42, §2.2] and [18, §4.1] for a thorough discussion  [^2] Note that the action of the special conformal transformations on a point in $\mathbb{R}^{p,q}$ can be interpreted as an inversion (about the origin), followed by a translation, followed by another inversion.
@@ 2.1.1. A more physical formulation
We use the customary notation from the physics literature where repeated indices are summed over, e.g., $\partial^{\mu}x_{\mu}=\sum_{\mu=1}^{d=p+q}\partial^{\mu}x_{\mu}$. Furthermore, indices are “raised” and “lowered” by contracting with the metric (tensor) $g_{\mu\nu}$ or its inverse. For example, $\partial_{\mu}=g_{\mu\nu}\partial^{\nu}=\sum_{\nu=1}^{d}g_{\mu\nu}\partial^{\nu}$. The metric is given by $g_{\mu\nu}=\text{diag}(\mathbb{1}_{p},-\mathbb{1}_{q})$.
@@ 2.2. The Superconformal Case
We will focus on superconformal Lie algebras, as they play the key role in this paper and because they are simpler to describe in technical terms (In particular, we will not discuss super analogs of the conformal compactification.) Our approach is borrowed from Shnider’s elegant and efficient discussion [43]. We begin by recalling how one formulates the super versions of the spacetimes $\mathbb{R}^{p,q}$ and of their isometries. That is, we review what super-Minkowski space is and what the super Poincaré group is (in arbitrary signature) to elucidate the logic that motivates a formulation of superconformal algebras.
@@
Remark 2.9.
In the physics literature it is typical for super Poincaré groups to be referred to by an index usually denoted by $\mathcal{N}$ combinatorially describing spinorial representations of $\mathrm{Spin}(p,q)$. If there is a single irreducible spinorial representation $S$, then the super Poincaré algebra associated to $S^{\oplus k}$ is called the $\mathcal{N}=k$ super Poincaré algebra and denoted $\mathrm{ISO}(p,q|k)$. If there are two irreducible spinorial representations $S_{+},S_{-}$, then the super Poincaré algebra associated to $S_{+}^{k_{+}}\oplus S_{-}^{k_{-}}$ is called the $\mathcal{N}=(k_{+},k_{-})$ super Poincaré algebra and denoted $\mathrm{ISO}(p,q|k_{+},k_{-})$.
@@
In looking for a super version of the conformal algebra, we want to construct a super Lie algebra $\mathfrak{sconf}(p,q|S)$ that depends on this choice of a Majorana spinor representation $S$ and the bilinear form. It should include the super Poincaré Lie algebra as a subalgebra, just as the Poincaré Lie algebra is a subalgebra of $\mathfrak{conf}(p,q)$. Moreover, the even part of $\mathfrak{sconf}(p,q|S)$ should contain the conformal algebra $\mathfrak{conf}(p,q)$ as a Lie subalgebra. In other words, $\mathfrak{sconf}(p,q|S)$ should be a common generalization of usual supersymmetry algebra and usual conformal algebra.
@@
When studying super Lie algebras $\mathfrak{g}$, we will use the notation [,][-,-] for the full graded skew-symmetric Lie bracket
$\colon\wedge^{2}(\mathfrak{g})\to\mathfrak{g}.$
@@
The remarkable theorem is that such a super Lie algebra exists only when $p+q\leq 6$, as discovered by Nahm [37], although we emphasize here the clean mathematical approach of Shnider [43], who articulates these properties as explicit hypotheses. Shnider begins by studying the complexification of the situation above, since any such real super Lie algebras has a complexifiction and there is no dependence on a choice of signature after complexifying. He then shows that if such a complex super Lie algebra $\mathfrak{g}$ exists, there is a simple super Lie algebra, satisfying the same properties, which is a quotient or subalgebra of $\mathfrak{g}$. Moreover, its bosonic part splits as the sum of $\mathfrak{so}(p+q,\mathbb{C})$ and a complementary ideal. He then uses Kac’s classification of simple super Lie algebras to show that such complex super Lie algebras do not exist above dimension 6. Identifying the relevant simple super Lie algebras in dimensions less than or equal to 6 is possible and already accomplished.
@@
As in Remark 2.9 the superconformal algebras are often indexed following the combinatorial indexing for the spinorial representation in their super Poincaré summand. Hence one refers to $\mathcal{N}=k$ superconformal algebras in dimension 3,4 and $\mathcal{N}=(k,0)$ superconformal algebras in dimension 6. We will use this terminology below.
Remark 2.16.
The classification of simple super Lie algebras has played an essential role in the physics literature on the structure and properties of superconformal theories [34, 36, 12]. In particular, Dolan and Osborn [19] developed the representation theory of 4d superconformal algebras, and Córdova, Dumitrescu and Intriligator [14] provided a thorough analysis of the representation theory of superconformal algebras in dimensions 2 to 6 and its applications to superconformal field theory.